A right triangle has a height of 9 yards and a hypotenuse of 12 yards, what is its area?
Answer by jim_thompson5910(35256)

The missing leg is sqrt(12^2 - 9^2) = sqrt(144-81) = sqrt(63) = 3*sqrt(7), which is the base.

So the exact area is

A = (b*h)/2

A = (3*sqrt(7)*9)/2

A = (27*sqrt(7))/2

which approximates to

A = 35.717642699 square yards.